# Sensor-Based Distributionally Robust Control for Safe Robot Navigation in Dynamic Environments

### v1

@@ 7.2 CLF-Based Path Following
(Practical considerations for control bounds): The original CLF-CBF QP formulation in (6) assumes no control bounds, a condition often not met in real-world robot applications due to physical limitations, such as maximum speed and acceleration. To ensure the applicability of our approach within these practical constraints, one can tune the parameters $k$ and $\zeta$ in the governor dynamics (24) to establish a smooth path-following behavior that respects the control bounds. $\bullet$
@@
8 CLF DR-CBF Formulation for Unicycle Dynamics
@@
Figure 4: (a) A robot is depicted following a path generated by a motion planning algorithm, and a dynamic local reference goal is highlighted in yellow. (b) The robot sensing the environment with a $360$-degree LiDAR sensor mounted at $\tilde{\mathbf{x}}$. The CBF samples $\{h_{i}(\mathbf{x})\}_{i=1}^{N}$ are the rays with boundary points highlighted as red triangles, and are selected based on the distance from the LiDAR detections to the robot body.
In this section, we instantiate the control strategy developed in the previous sections to the case of unicycle dynamics in (7). We first design a CLF that enables stabilization to a desired goal point. Then, we discuss the validity of using a signed distance function (SDF) (Hoppe et al., 1992) as a CBF candidate for ensuring safety. Furthermore, we provide a discussion on how to select CBF samples based on range sensor measurements, which is crucial for the practical implementation of the distributionally robust safety constraint. By combining the unicycle-specific CLF and the data-driven CBF, we obtain a CLF-DR-CBF QP formulation tailored to the unicycle dynamics, enabling safe and efficient navigation in unknown dynamic environments.
@@ 8.1 Unicycle Stabilization to a Goal Position
We design a CLF that enables stabilization of the unicycle dynamics (7) to a desired goal point $\mathbf{q}_{G}\in\mathbb{R}^{2}$. We define the state space of the unicycle as ${\cal X}={\cal D}\times[-\pi,\pi)$, where ${\cal D}\subset\mathbb{R}^{2}$ is a sufficiently large compact set containing the environment of interest, including the goal point $\mathbf{q}_{G}$ and the planned path $\gamma$. This ensures that ${\cal X}$ is compact, satisfying Assumption 7.1.
@@ 8.2 Signed Distance Function as Control Barrier Function
In unknown and dynamic environments, the precise computation of the barrier function $h(\mathbf{x},t)$ or the construction of a probabilistic model is challenging. Therefore, this section focuses on safe control synthesis, directly leveraging the distance measurements $\boldsymbol{\eta}(\phi(\mathbf{x})_{s},\mathbf{R}_{s})=[\eta_{1}(\phi(\mathbf{x})_{s},\mathbf{R}_{s}),...,\eta_{K}(\phi(\mathbf{x})_{s},\mathbf{R}_{s})]^{\top}$ to define the safety constraint.
Given that the range sensor is mounted at $\phi(\mathbf{x})_{s}=[x+l\cos(\theta),y+l\sin(\theta)]^{\top}\in\mathbb{R}^{2}$, we describe the dynamics of $\mathbf{x}_{s}=[x+l\cos(\theta),y+l\sin(\theta),\theta]^{\top}$ as outlined by Cortés and Egerstedt (2017):
$\dot{\mathbf{x}}_{s}=\underbrace{\left[\begin{matrix}\cos(\theta)&-l\sin(\theta)\\ \sin(\theta)&l\cos(\theta)\\ 0&1\end{matrix}\right]}_{\tilde{\mathbf{g}}(\mathbf{x})}\left[\begin{matrix}v\\ \omega\end{matrix}\right].$ (34)
Next, we discuss the validity of using the signed distance function from the robot position to the obstacle set ${\cal O}(t)$ as a time-varying CBF candidate for the dynamics (34). The signed distance function $\varphi:\mathbb{R}^{2}\times\mathbb{R}\mapsto\mathbb{R}$ of set ${\cal O}(t)$ is:
$\varphi(\phi(\mathbf{x}),t):=\begin{cases}-d(\phi(\mathbf{x}),\partial{\cal O}(t)),&\phi(\mathbf{x})\in{\cal O}(t),\\ \phantom{-}d(\phi(\mathbf{x}),\partial{\cal O}(t)),&\phi(\mathbf{x})\notin{\cal O}(t).\end{cases}$ (35)
The distance function provides a useful measure of safety, with positive values indicating free space. We exploit this property to construct a CBF that ensures the safety of the system dynamics described in (34).
Let $r_{0}$ denote the robot radius, as shown in Fig. 4(b). To ensure the safety of the entire robot body ${\cal B}(\mathbf{x})$, we define the candidate CBF as
$h(\mathbf{x},t)=\varphi(\phi(\mathbf{x})_{s},t)-r_{0}-l,$ (36)
which ensures that the safety margin accounts for both the robot body and the placement of the LiDAR. The next result shows that $h(\mathbf{x},t)$ in (36) is a valid time-varying CBF candidate for the dynamics in (34).
@@
(Signed distance function as a time-varying control barrier function): Assume that the partial time derivative of the signed distance function is bounded, i.e., there exists a constant $B>0$ such that $\left|\frac{\partial\varphi(\phi(\mathbf{x})_{s},t)}{\partial t}\right|\leq B$ for all $\phi(\mathbf{x})_{s}\in{\cal D}\subset\mathbb{R}^{2}$ and $t\in\mathbb{R}$. Then, the function $h(\mathbf{x},t)$ in (36) is a valid time-varying CBF for the system dynamics (34).
@@
From the dynamics model (34), the derivative of the candidate time-varying CBF along the system dynamics is:
$\dot{h}$ $(\mathbf{x},t)=\nabla\varphi(\phi(\mathbf{x})_{s},t)^{\top}\dot{\phi(\mathbf{x})}_{s}+\frac{\partial\varphi(\phi(\mathbf{x})_{s},t)}{\partial t}$
$=\nabla\varphi(\phi(\mathbf{x})_{s},t)^{\top}\begin{bmatrix}v\cos(\theta)\!-\!l\omega\sin(\theta)\\ v\sin(\theta)\!+\!l\omega\cos(\theta)\end{bmatrix}\!+\!\frac{\partial\varphi(\phi(\mathbf{x})_{s},t)}{\partial t}$
$=\nabla\varphi(\phi(\mathbf{x})_{s},t)^{\top}R(\theta)\mathbf{u}_{l}+\frac{\partial\varphi(\phi(\mathbf{x})_{s},t)}{\partial t},$
where $\mathbf{u}_{l}=[v,l\omega]^{\top}$, $R(\theta)=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}$. It is known (Crandall and Lions, 1983) that the gradient $\nabla\varphi(\phi(\mathbf{x})_{s},t)$ of the signed distance function is well-defined and has a unit norm near the boundary of the obstacle set ${\cal O}(t)$ Next, by choosing $\mathbf{u}_{l}=R(-\theta)\mathbf{p}$, where $\mathbf{p}=\beta\nabla\varphi(\phi(\mathbf{x})_{s},t)$, $\beta>0$ is a scaling factor, we have
$\dot{h}(\mathbf{x})=\beta\|\nabla\varphi(\phi(\mathbf{x})_{s},t)\|^{2}+\frac{\partial\varphi(\phi(\mathbf{x})_{s},t)}{\partial t}\geq\beta-B.$ (37)
The scaling factor $\beta$ can be chosen sufficiently large such that $\beta>B$, ensuring $\dot{h}(\mathbf{x},t)>0$ near the boundary of the obstacle set. Therefore, $h(\mathbf{x},t)$ in (36) is a valid time-varying CBF for (34). ∎∎
@@
From a practical perspective, it is important to note that, if there are bounds on the control inputs, e.g., $|v|\leq v_{\max}$ and $|\omega|\leq\omega_{\max}$, then the scaling factor $\beta$ in the proof of Lemma 8.2 must be chosen to ensure that $\mathbf{u}_{l}$ satisfies these constraints while maintaining $\dot{h}(\mathbf{x},t)>0$ near the boundary of the obstacle set. This means that the bound $B$ on the partial time derivative of the distance function must be compatible with the control bounds to ensure the inequality (37) holds. In other words, the obstacles must move at speeds compatible with the motion capabilities of the unicycle to guarantee safety.

### v2

@@ 7.2 CLF-Based Path Following
In practical implementations, the control input is applied at a finite sampling rate. To preserve stability guarantees, we use a sampling frequency of 50 Hz in the simulations and experiments, ensuring that the discretized behavior closely approximates the continuous-time result.
@@
(Practical considerations for control bounds): The original CLF-CBF QP formulation in (6) assumes no control bounds, a condition often not met in real-world robot applications due to physical limitations, such as maximum speed and acceleration. To ensure the applicability of our approach within these practical constraints, one can tune the parameters $k$ in the governor dynamics (27) to establish a smooth path-following behavior that respects the control bounds. Specifically, increasing $k$ slows down the progression of the reference goal $\gamma(g)$, reducing the control effort required. $\bullet$
@@
8 Application to Differential-Drive Robot
@@
Figure 4: (a) A robot is depicted following a path generated by a motion planning algorithm, and a dynamic local reference goal is highlighted in yellow. (b) The robot senses the environment with a $360$-degree LiDAR sensor mounted at $\tilde{\mathbf{x}}$. The CBF samples $\{h_{i}(\mathbf{x})\}_{i=1}^{N}$ are the rays with boundary points highlighted as red triangles, and are selected based on the distance from the LiDAR detections to the robot body.
While the theory of our control method is applicable to general control-affine systems (1), our experiments focus on a wheeled differential-drive robot.
We consider a robot with state $\mathbf{x}:=[x,y,\theta]^{\top}\in{\cal X}\subseteq\mathbb{R}^{2}\times[-\pi,\pi)$, input $\mathbf{u}:=[v,\omega]^{\top}\in{\cal U}\subset\mathbb{R}^{2}$, and dynamics:
The function $\phi:{\cal X}\rightarrow\mathbb{R}^{2}$ projecting the robot state $\mathbf{x}$ to its position is $\phi(\mathbf{x})=\phi([x,y,\theta]^{\top})=[x,y]^{\top}$.
We define the state space of the robot as ${\cal X}={\cal D}\times[-\pi,\pi)$, where ${\cal D}\subset\mathbb{R}^{2}$ is a sufficiently large compact set containing the environment of interest, including the goal point $\mathbf{q}_{G}$ and the planned path $\gamma$.
In this section, we instantiate the control strategy developed in the previous sections to the case of differential-drive dynamics in (32). We first design a CLF that enables stabilization to a desired goal point, as requested by Assumption 7.2. Then, we discuss the validity of using the SDF as a CBF candidate for ensuring safety. Next, we provide a discussion on how to select CBF samples based on range sensor measurements, which is crucial for the practical implementation of the proposed DR-CBF formulation. To address uncertainties in both localization and sensor data, we also introduce a unified error model that captures how localization errors propagate to LiDAR measurements, impacting the robot’s perception of obstacles. By combining the unicycle-specific CLF and the data-driven CBF, we obtain a CLF-DR-CBF QP formulation tailored to the unicycle dynamics, enabling safe and efficient navigation in unknown dynamic environments.
@@ 8.1 Unicycle Stabilization to a Goal Position
We design a CLF that enables stabilization of the unicycle dynamics (32) to a desired goal point $\mathbf{q}_{G}\in\mathbb{R}^{2}$.
@@ 8.2 Signed Distance Function as Control Barrier Function
In unknown and dynamic environments, the precise computation of the barrier function $h(\mathbf{x},t)$ or the construction of a probabilistic model is challenging. This section focuses on showing that the robot SDF (9) is a valid CBF for (32) under appropriate assumptions.
We make the following assumptions:
Assumption 8.2.
(Regularity of Robot Shape): The robot shape ${\cal B}_{0}$ has a smooth boundary and is compact.
Based on the robot SDF in (9), we define the candidate CBF as:
$h(\mathbf{x},t)\!=\!d({\cal B}(\mathbf{x}),{\cal O}(t))\!:=\!\inf_{\mathbf{q}\in{\cal O}(t)}d_{{\cal B}_{0}}\left(\mathbf{R}(\mathbf{x})^{\top}(\mathbf{q}-\phi(\mathbf{x}))\right),$ (38)
where ${\cal O}(t)\subset\mathbb{R}^{2}$ is the obstacle set at time $t$, and $\mathbf{R}(\mathbf{x})\in\text{SO}(2)$ is the rotation matrix that describes the unicycle’s orientation. By Assumption 8.2, we know that $d_{{\cal B}_{0}}(\cdot)$ is 1-Lipschitz continuous (Fitzpatrick, 1980). The presence of the infimum operator in (38) ensures that the candidate CBF $h(\mathbf{x},t)$ is Lipschitz continuous. However, it is important to note that $h(\mathbf{x},t)$ is, in general, not smooth due to both the properties of $d_{{\cal B}_{0}}(\cdot)$ and the infimum operation.
Therefore, we review some concepts from nonsmooth analysis to handle the potential non-differentiability of $h$.
Definition 8.3.
(Generalized Gradient): Let $h:{\cal X}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ be Lipschitz continuous near $\mathbf{z}:=(\mathbf{x},t)\in\mathbb{R}^{4}$, and suppose $\mathcal{Z}$ is any set of Lebesgue measure zero in $\mathbb{R}^{4}$. The Clarke generalized gradient (Clarke, 1981) of $h$ at $\mathbf{z}$ is defined as
$\partial h(\mathbf{z})=\operatorname{co}\left\{\lim_{i\to\infty}\nabla h(\mathbf{z}_{i})\,\big{|}\,\mathbf{z}_{i}\to\mathbf{z},\,\mathbf{z}_{i}\notin\Omega_{h}\cup\mathcal{Z}\right\},$ (39)
where $\operatorname{co}$ denotes the convex hull and $\Omega_{h}$ is the set of points where $h$ fails to be differentiable.
Using this definition, we extend the notion of time-varying control barrier functions to functions that are not necessarily differentiable everywhere.
Definition 8.4.
(Nonsmooth Time-Varying Control Barrier Function): A Lipschitz continuous and regular function $h:{\cal X}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ is called a nonsmooth time-varying control barrier function on a set $\mathcal{H}\subset\mathbb{R}^{4}$ if there exists an extended class-$\mathcal{K}$ function $\alpha$ such that for all $\mathbf{x}\in{\cal X}$ and all $t\geq 0$,
$\sup_{\mathbf{u}\in\mathcal{U}}\inf_{(\mathbf{y},z)\in\partial h(\mathbf{x},t)}\left\{\mathbf{y}^{\top}\left(f(\mathbf{x})+g(\mathbf{x})\mathbf{u}\right)+z\right\}\geq-\alpha(h(\mathbf{x},t)),$ (40)
where $\partial h(\mathbf{x},t)$ is the generalized gradient of $h$.
Next, we show that $h(\mathbf{x},t)$ in (38) is a valid time-varying CBF candidate for the dynamics in (32) under the following assumptions.
Assumption 8.5.
(Boundedness and Regularity of Obstacles): We assume that:
  • The initial obstacle set $\mathcal{O}_{0}$ is bounded, and the transformations $A_{\mathcal{O}}(t)$ and $\mathbf{b}_{\mathcal{O}}(t)$ in (10) are uniformly bounded over $t$.
  • There exists a constant $B>0$ such that for all $(\mathbf{x},t)$ and all $(\mathbf{y},z)\in\partial h(\mathbf{x},t)$, it holds that $|z|\leq B$.
The following result establishes the boundedness of the obstacle set ${\cal O}(t)$. Its proof is straightforward and therefore omitted.
@@
(Obstacle Set Boundedness): Under Assumption 8.5, the obstacle set ${\cal O}(t)$ is bounded for all $t\geq 0$.
Next, we show that $h(\mathbf{x},t)$ is Lipschitz.
Lemma 8.7.
(Lipschitzness of Non-smooth CBF): Under Assumptions 8.2 and 8.5, the function $h(\mathbf{x},t)$ is Lipschitz in $(\mathbf{x},t)$.
@@
For each fixed $\mathbf{q}\in\mathcal{O}(t)$, we write
$h_{\mathbf{q}}(\mathbf{x})=d_{\mathcal{B}_{0}}\left(\mathbf{R}(\mathbf{x})^{\top}\left(\mathbf{q}-\phi(\mathbf{x})\right)\right).$
Since $\mathbf{R}(\mathbf{x})$ and $\phi(\mathbf{x})$ are continuously differentiable, they are also Lipschitz continuous on the compact set ${\cal X}$. Let $L_{\mathbf{R}}$ and $L_{\phi}$ be their Lipschitz constants, respectively, so that $\left\|\mathbf{R}(\mathbf{x}_{1})-\mathbf{R}(\mathbf{x}_{2})\right\|\leq L_{\mathbf{R}}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$ and $\left\|\phi(\mathbf{x}_{1})-\phi(\mathbf{x}_{2})\right\|\leq L_{\phi}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$. Now, for any $\mathbf{x}_{1},\mathbf{x}_{2}\in{\cal X}$ and fixed $\mathbf{q}\in\mathcal{O}(t)$, one has
$\left|h_{\mathbf{q}}(\mathbf{x}_{1})-h_{\mathbf{q}}(\mathbf{x}_{2})\right|$
$=\left|d_{\mathcal{B}_{0}}\left(\mathbf{R}(\mathbf{x}_{1})^{\top}\left(\mathbf{q}\!-\!\phi(\mathbf{x}_{1})\right)\right)\!-\!d_{\mathcal{B}_{0}}\left(\mathbf{R}(\mathbf{x}_{2})^{\top}\left(\mathbf{q}\!-\!\phi(\mathbf{x}_{2})\right)\right)\right|$
$\leq\left\|\mathbf{R}(\mathbf{x}_{1})^{\top}\left(\mathbf{q}-\phi(\mathbf{x}_{1})\right)-\mathbf{R}(\mathbf{x}_{2})^{\top}\left(\mathbf{q}-\phi(\mathbf{x}_{2})\right)\right\|$
$\leq\left\|\left[\mathbf{R}(\mathbf{x}_{1})^{\top}-\mathbf{R}(\mathbf{x}_{2})^{\top}\right]\left(\mathbf{q}-\phi(\mathbf{x}_{1})\right)\right\|+$
$\quad\left\|\mathbf{R}(\mathbf{x}_{2})^{\top}\left(\phi(\mathbf{x}_{2})-\phi(\mathbf{x}_{1})\right)\right\|.$ (41)
where the first equality comes from the fact that $d_{{\cal B}_{0}}$ is 1-Lipschitz (Fitzpatrick, 1980).
For the first term in (41), we have
$\left\|\left[\mathbf{R}(\mathbf{x}_{1})^{\top}-\mathbf{R}(\mathbf{x}_{2})^{\top}\right]\left(\mathbf{q}-\phi(\mathbf{x}_{1})\right)\right\|\\ \leq L_{\mathbf{R}}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\cdot\left(\left\|\mathbf{q}\right\|+\left\|\phi(\mathbf{x}_{1})\right\|\right).$
Since ${\cal X}$ is bounded by assumption and $\mathcal{O}(t)$ is also bounded, cf. Lemma 8.6, there exist $M_{\mathbf{q}},M_{\phi}>0$ such that $\left\|\mathbf{q}\right\|\leq M_{\mathbf{q}}$ and $\left\|\phi(\mathbf{x}_{1})\right\|\leq M_{\phi}$. Therefore:
$\left\|\left[\mathbf{R}(\mathbf{x}_{1})^{\top}-\mathbf{R}(\mathbf{x}_{2})^{\top}\right]\left(\mathbf{q}-\phi(\mathbf{x}_{1})\right)\right\|\\ \leq L_{\mathbf{R}}(M_{\mathbf{q}}+M_{\phi})\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$
For the second term in (41), since $\left\|\mathbf{R}(\mathbf{x})\right\|=1$, we have
$\left\|\mathbf{R}(\mathbf{x}_{2})^{\top}\left(\phi(\mathbf{x}_{2})-\phi(\mathbf{x}_{1})\right)\right\|\leq L_{\phi}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$
Combining the bounds from the two terms, we conclude
$\left|h_{\mathbf{q}}(\mathbf{x}_{1})-h_{\mathbf{q}}(\mathbf{x}_{2})\right|\leq L_{h}\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|,$ (42)
with $L_{h}=(L_{\mathbf{R}}(M_{\mathbf{q}}+M_{\phi})+L_{\phi})$. As established in Rockafellar and Wets (2009, Proposition 9.10), the pointwise infimum of a collection of Lipschitz functions is Lipschitz with a constant bounded by the maximum of the individual Lipschitz constants. Therefore, since $h_{\mathbf{q}}(\mathbf{x})$ is uniformly Lipschitz with constant $L_{h}$ and $\mathcal{O}(t)$ is bounded, the infimum $h(\mathbf{x},t)=\inf_{\mathbf{q}\in\mathcal{O}(t)}h_{\mathbf{q}}(\mathbf{x})$ is Lipschitz with the same constant $L_{h}$.
Furthermore, by Assumption 8.5, $h(\mathbf{x},t)$ is Lipschitz in $t$ with constant $B$. ∎∎
Next, our goal is to show that for all $(\mathbf{x},t)\in{\cal X}\times\mathbb{R}_{\geq 0}$, there exists a control input $\mathbf{u}\in\mathcal{U}$ such that the CBF condition (40) holds.
Proposition 8.8.
(Sufficient Conditions for Signed Distance Function as a Nonsmooth Time-Varying Control Barrier Function): Under Assumptions 8.2 and 8.5, the function $h(\mathbf{x},t)$ defined in (38) satisfies the nonsmooth CBF condition (40) for the unicycle dynamics (32), provided that $\inf_{(\mathbf{0},z)\in\partial h(\mathbf{x},t)}z\geq-\alpha_{h}(h(\mathbf{x},t))$.
Proof.
Using the unicycle dynamics (32), for any $(\mathbf{y},z)\in\partial h(\mathbf{x},t)$, the expression inside the infimum in (40) becomes
$\mathbf{y}^{\top}\left(f(\mathbf{x})+g(\mathbf{x})\mathbf{u}\right)+z=$ (43)
$\mathbf{y}^{\top}\left(\begin{bmatrix}\cos\theta\\ \sin\theta\\ 0\end{bmatrix}v+\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}\omega\right)+z=a_{v}v+a_{\omega}\omega+z,$
where
$a_{v}$ $=\mathbf{y}^{\top}\begin{bmatrix}\cos\theta\\ \sin\theta\\ 0\end{bmatrix},\quad a_{\omega}=\mathbf{y}^{\top}\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}.$
From Lemma 8.7, $\|y\|\leq L_{h}$. It follows that $|a_{v}|,\;|a_{\omega}|\leq L_{h}$. We now consider two cases:
Case 1: $(\boldsymbol{0},z)\notin\partial h(\mathbf{x},t)$. In this case, $\mathbf{y}\neq\boldsymbol{0}$, so at least one of $a_{v}$ or $a_{\omega}$ is non-zero. Define $\varepsilon=a_{v}^{2}+a_{\omega}^{2}>0$ and choose
$v=\frac{a_{v}}{\varepsilon}\beta,\quad\omega=\frac{a_{\omega}}{\varepsilon}\beta,$ (44)
where $\beta>0$ is a scaling factor. Then, since $z\in[-B,B]$ (cf. Assumption 8.5), we get
$a_{v}v+a_{\omega}\omega+z$ $=\left(\frac{a_{v}^{2}+a_{\omega}^{2}}{\varepsilon}\right)\beta+z\geq\beta-B.$ (45)
The scaling factor $\beta$ can be chosen sufficiently large to ensure that the nonsmooth CBF condition (40) is satisfied.
Case 2: $(\boldsymbol{0},z)\in\partial h(\mathbf{x},t)$. In this case, the spatial gradient of $h$ vanishes, and the expression in (43) simplifies to:
$\mathbf{y}^{\top}\left(f(\mathbf{x})+g(\mathbf{x})\mathbf{u}\right)+z=z.$
The nonsmooth CBF condition (40) holds in this case because of the hypotheses of the statement. ∎∎
Regarding the requirement of Proposition 8.8, intuitively, when $\mathbf{y}=\boldsymbol{0}$, the spatial gradient does not provide directional guidance for the robot, and the safety condition depends solely on the temporal evolution of $h$, i.e., the behavior of the obstacles. The requirement $\inf_{(\mathbf{0},z)\in\partial h(\mathbf{x},t)}z\geq-\alpha_{h}(h(\mathbf{x},t))$ ensures that the nonsmooth CBF condition is satisfied in this case.
@@
In practice, we need to ensure that the control inputs $v$ and $\omega$ satisfy the control constraints $|v|\leq v_{\max}$ and $|\omega|\leq\omega_{\max}$. Based on (44) and (45), this means that the bound $B$ on the partial time derivative of the distance function must be compatible with the control bounds to ensure the inequality (40) holds. In other words, to guarantee safety, the obstacles must move at speeds compatible with the motion capabilities of the robot.